OpenStudy (anonymous):

Tell why a number divisible by 9 is also divisible by 3

OpenStudy (anonymous):

It helps if you do some examples. 4x9=36 This means, by definition, that 36 is divisible by 9. The next step would be dividing 36 by 3. 36/3=12. Try multiplying any number by 9 and then dividing that answer by three. If you can successfuly divide this by three and get a whole number, then the number is divisible by 3 as well as 9.

OpenStudy (anonymous):

because 3 is a multiple of 9 I will teach you a little more if you take any number and add the componets and take there sum and if its multiple of three is also is dividisible by three example 63 then 6+3=9 and 3 goes into nine so 63 is divisible by three exaple two 123 is 1+2+3=6 so 123 is divisible by 3. get it?

OpenStudy (unklerhaukus):

\[b|a \quad \top\]\[\implies a=q\times b+r,\quad q\in \mathbb Z,r=0\] \[3|9 \quad \top\]\[\because 9=3\times3+0\]

OpenStudy (anonymous):

Another way of looking at it is be "taking it apart." If we use the same example above we can see this. 4x9=36, as we have already determined. Next, we should figure out the factors of the two factors in this equation, 2 and 9. The factors of 4 we can use are 2, the factors of 9 we can use are 3. (2x2) x (3x3)=36 If this is true, again by definition, 36 is divisible by both 9 and 3.